if f (x) = 5x - 1, find the following:
f (a - 2)

Answer by funmath(2933)

if f (x) = 5x - 1, find the following:
f (a - 2)
Substitute a-2 in where x is in the function.
f(a-2)=5(a-2)-1
f(a-2)=5a-10-1
f(a-2)=5a-11
Happy Calculating!!!